OpenStudy (anonymous):

Find the sum of the geometric sequence. (3/2),(3/8),(3/32),(3/128),(3/512) PLEASE HELP ! I GIVE MEDALS

OpenStudy (anonymous):

Just add up the fractions... Or terms.

OpenStudy (anonymous):

THats not the answer

OpenStudy (anonymous):

@amoodarya and ideas ?

OpenStudy (amoodarya):

s=3/2+3/8+3/32+3/128 +3/512 +... miltiply by 1/4 (r=1/4) 1/4s= 3/8+3/32+3/128 +3/512 +... now s-1/4s=3/2 so 3/4s=3/2

OpenStudy (anonymous):

I dont see that ..... these are the options @amoodarya

OpenStudy (anonymous):

What did I say? Add up the fractions! That's the sum! the final result after adding is the sum! So the answer is C.

OpenStudy (anonymous):

what about Write the sum using summation notation, assuming the suggested pattern continues. 16 + 25 + 36 + 49 + ... + n2 + ... @Zelda

OpenStudy (amoodarya):

1^2+2^2+3^2+.....n^2=n(n+1)(2n+1)/6 so 4^2+5^2+6^2+....n^2=1^2+2^2+3^2+.....n^2-(1+4+9)=n(n+1)(2n+1)/6-(1+4+9)

OpenStudy (anonymous):

so which option ? @amoodarya

OpenStudy (anonymous):

and for which question ?

OpenStudy (amoodarya):

s=3/2+3/8+3/32+3/128 +3/512 s=3/512(256+64+16+4+1) s=1023/512

OpenStudy (anonymous):

can you help me with that second one i just commented

OpenStudy (anonymous):

@amoodarya

OpenStudy (amoodarya):

second question did you mean 4^2+5^2+...n^2= ? or 4^2+5^2+...n^2+....= ?

OpenStudy (anonymous):

the second one @amoodarya

OpenStudy (amoodarya):

yeah the second question

OpenStudy (anonymous):

4^2+5^2+...n^2+....= @amoodarya

OpenStudy (anonymous):

for this question 16 + 25 + 36 + 49 + ... + n2 + ...

OpenStudy (amoodarya):

1^2+2^2+3^2+.....n^2+... =infinity but 1^2+2^2+3^2+.....n^2=n(n+1)(2n+1)/6 so 4^2+5^2+6^2+....n^2=1^2+2^2+3^2+.....n^2- (1+4+9)=n(n+1)(2n+1)/6- (1+4+9)

OpenStudy (anonymous):

The options are Im just having a hard time reading that

OpenStudy (anonymous):

@amoodarya

OpenStudy (anonymous):

D.

OpenStudy (amoodarya):